# Hölder regularity of Siciak-Zaharjuta extremal functions on compact Hermitian manifolds

### 05 Apr 2026 v2

@@ Abstract.
For a compact subset in a compact Hermitian manifold, we prove that the Hölder continuity of the extremal function at a given point in the set is a local property and that the Hölder continuity of a weighted extremal function follows from the Hölder continuities of the extremal function and the weight function with a uniform density in capacity. The second result can be seen as an extension of a result of Lu, Phung and Tô [LPT21]. Moreover, for a compact subset in a compact Hermitian manifold, we prove that the Hölder continuity of the extremal function with the uniform density in capacity is equivalent to the local Hölder continuity property, which is also equivalent to the weak local Hölder continuity property. These results are generalizations of the results of Nguyen [Ng24] on compact Kähler manifolds. We also show that the $\mu$-Hölder continuity property of a convex compact subset in $\mathbb{C}^{n}$ implies the local $\mu$-Hölder continuity property of order $1$.
@@ 1. Introduction
(1.2) $V_{E,\phi}(z)=V_{\omega;E,\phi}(z):=\sup\{v(z):v\in PSH(X,\omega),\;v|_{E}\leq\phi\}.$
When the domain of a real-valued function $\psi$ contains $E$, we write $V_{E,\psi|_{E}}$ as $V_{E,\psi}$ for convenience. There is a natural one-to-one correspondence between the Lelong class $\mathcal{L}$ and $PSH(\mathbb{CP}^{n},\omega_{FS})$ ([GZ17, Chapter 8]). When $\phi$ is a continuous function on a compact subset of $X$, we can extend $\phi$ to a continuous function on $X$ with the same supremum norm by Tietze’s extension theorem. An extension of a Hölder continuous function from a compact subset to $X$ is also possible with the same Hölder exponent and the same Hölder constant by Mcshane’s extension [Mc34].
@@
For $\mathbb{K}=\mathbb{R}$ or $\mathbb{K}=\mathbb{C}$, the closure of a bounded Lipschitz domain in $\mathbb{K}^{n}$ and the closure of a bounded domain in $\mathbb{K}^{n}$ with the geometrical condition, explained in Definition 1.8, are shown to have the local Hölder continuity property as subsets of $\mathbb{C}^{n}$ in [Ng24, Example 4.5]. In particular, the real cube $0,1]^{n}$, the polydisk $\bar{\mathbb{B}}(\mathbf{0},1)^{n}$ and a convex compact subset of $\mathbb{K}^{n}$ with the nonempty interior in $\mathbb{K}^{n}$ has a uniform density in capacity in $\mathbb{C}^{n}$. In [Ng24, Example 4.5], the uniform interior sphere condition implies the geometrical condition since the union of uniform closed balls conjecture was proved by Nour, Stern and Takche [NST11]. [Ng24, Lemma 4.7-(a)] directly showed that the closure of a bounded domain in $\mathbb{C}^{n}$ with a uniform interior sphere condition implies the local Hölder continuity property with improvement in both the exponent and the order compared to the geometric condition. This direct proof is also valid since the uniform interior $r$-sphere condition implies the union of uniform closed $r/2$-balls by [NST11].
@@
$\alpha(K):=\frac{2\mu}{q+2}\quad\text{and}\quad\alpha^{\prime}(K):=\frac{4\mu^{2}}{(q+2)(2\mu+(q+2)^{2})}$
@@
The Hölder continuity property of a general compact subset in $\mathbb{C}^{n}$, introduced in Section 3, does not imply the local Hölder continuity property, introduced in Section 3, of the subset. However, if a compact subset in $\mathbb{C}^{n}$ is convex, then the author proves that the Hölder continuity property of the subset implies the local Hölder continuity property of it with the same exponent and the order $1$.
@@
Let $F$ be a convex compact subset of $\mathbb{C}^{n}$ and $\mu\in(0,1]$. If $F$ has $\mu$-Hölder continuity property, then $F$ has local $\mu$-Hölder continuity property of order $1$.
@@
Let $\mathbb{K}$ be either $\mathbb{R}$ or $\mathbb{C}$. Then a convex compact subset of $\mathbb{K}^{n}$ with nonempty interior in $\mathbb{K}^{n}$ has local $1/2$-Hölder continuity property of order $1$.
@@ 2. Basic properties of extremal functions
The basic properties introduced in this section are the analogues for the extremal functions of the compact sets in $\mathbb{C}^{n}$. For the details of these basic properties, we refer to [Ah26]. Note that for subsets $E$ and $O$ of $X$ (resp. of $\mathbb{C}^{n}$), if $O$ is open in $X$ (resp. in $\mathbb{C}^{n}$), then $\sup_{U}V_{E}=\sup_{U}V_{E}^{*}$ (resp. $\sup_{O}L_{E}=\sup_{O}L_{E}^{*}$).
@@ 3.2. Local Hölder continuity property
By the HCP condition, we have
@@
$L_{F}(z)\leq 2C\max\{d_{\mathbb{C}^{n}}(z,F),(d_{\mathbb{C}^{n}}(z,F))^{\mu}\},\quad z\in\mathbb{C}^{n}.$
Since $F$ is convex, for each $a\in F$ and $0<r\leq 1$, the affinely transformed set a+(Fa)/ra+(F-a)/r contains F. Since $F$ is compact, there exists a small constant $1>\epsilon>0$ such that $\bar{\mathbb{B}}(a,1/\epsilon)$ contains $F$ for each $a\in F$.
For each $a\in F$ and $0<r\leq 1$, let $T_{a,r\epsilon}:\mathbb{C}^{n}\to\mathbb{C}^{n}$ be the invertible affine transformation given by $T_{a,r\epsilon}(z):=(z-a)/(r\epsilon)+a$. Then we have
$L_{F\cap\bar{\mathbb{B}}(a,r)}(z)=L_{T_{a,r\epsilon}(F)\cap\bar{\mathbb{B}}(a,\frac{1}{\epsilon})}(T_{a,r\epsilon}(z))\leq L_{F}(T_{a,r\epsilon}(z)),\quad z\in\mathbb{C}^{n}.$
Now, we can show the local $\mu$-HCP of order $1$ of $F$. For each $z\in\mathbb{C}^{n}$, there exists $a_{z}\in F$ such that $d_{\mathbb{C}^{n}}(z,F)=\|z-a_{z}\|$. Then we get
$L_{F\cap\bar{\mathbb{B}}(a,r)}(z)$ $\leq L_{F}(T_{a_{z},r\epsilon}(z))$
$\leq 2C\max\{d_{\mathbb{C}^{n}}(T_{a_{z},r\epsilon}(z),F),(d_{\mathbb{C}^{n}}(T_{a_{z},r\epsilon}(z),F))^{\mu}\}$
$\leq 2C\max\{\frac{\|z-a_{z}\|}{r\epsilon},(\frac{\|z-a_{z}\|}{r\epsilon})^{\mu}\}$
$\leq\frac{2C}{\epsilon}\frac{\|z-a_{z}\|^{\mu}}{r},\quad d_{\mathbb{C}^{n}}(z,F)=\|z-a_{z}\|\leq 1.$
Therefore, $F$ has local $\mu$-HCP of order $1$. ∎
@@ 4.1. Demailly estimates
(4.3), (4.5) and [GKZ08, Remark 4.2](or [Z21, Theorem 3.2]) together give
@@ 5.3. Proof of the third Theorem
[Mc34], Mcshane extension, extends a Hölder continuous function from a compact subset to $X$ with the same Hölder exponent and the same Hölder constant. Since $X$ is compact, each connected component of $X$ has a finite diameter with respect to the Riemannian distance $d_{\omega}$. Therefore, for a given bounded subset $\{\phi_{\gamma}\}_{\gamma\in\Gamma}$ in $C^{\mu^{\prime}}(K)$, each $\phi_{\gamma}$ can be extended to a function $\tilde{\phi}_{\gamma}$ in $C^{\mu^{\prime}}(X)$ such that $\{\tilde{\phi}_{\gamma}\}_{\gamma\in\Gamma}$ is bounded in $C^{\alpha}(X)$. Let $\mu^{\prime\prime}:=(\mu^{\prime})^{2}/(\mu^{\prime}+2+q)$. The direct computation gives
$\mu^{\prime\prime}=\frac{4\mu^{2}}{(q+2)(2\mu+(q+2)^{2})}.$

### 06 May 2026 v3

@@ Abstract.
For a compact subset in a compact Hermitian manifold, we prove that the Hölder continuity of the extremal function at a given point in the set is a local property and that the Hölder continuity of a weighted extremal function follows from the Hölder continuities of the extremal function and the weight function with a uniform density in capacity. The second result can be seen as an extension of a result of Lu, Phung and Tô [LPT21]. Moreover, for a compact subset in a compact Hermitian manifold, we prove that, both at the point level and at the global level, the Hölder continuity of the extremal function with the uniform density in capacity is equivalent to the local Hölder continuity property, which is also equivalent to the weak local Hölder continuity property. These results are generalizations of the results of Nguyen [Ng24] on compact Kähler manifolds. We also show that the $\mu$-Hölder continuity property of a convex compact subset in $\mathbb{C}^{n}$ and of a compact subset in $\mathbb{C}^{n}$ at a star center imply the local $\mu$-Hölder continuity property of order $\mu$ of the convex compact subset and of the compact subset at the point, respectively.
@@ 1. Introduction
(1.2) $V_{E,\phi}(x)=V_{\omega;E,\phi}(x):=\sup\{v(x):v\in PSH(X,\omega),\;v|_{E}\leq\phi\},\quad x\in X.$
When the domain of a real-valued function $\psi$ contains $E$, we write $V_{E,\psi|_{E}}$ as $V_{E,\psi}$ for convenience. There is a bijection between the Lelong class $\mathcal{L}$ and $PSH(\mathbb{CP}^{n},\omega_{FS})$ ([GZ17, Chapter 8]). When $\phi$ is a continuous function on a compact subset of $X$, we can extend $\phi$ to a continuous function on $X$ with the same supremum norm by Tietze’s extension theorem. An extension of a Hölder continuous function from a compact subset to $X$ is also possible with the same Hölder exponent and the same Hölder constant by Mcshane’s extension [Mc34].
@@
For $\mathbb{K}=\mathbb{R}$ or $\mathbb{K}=\mathbb{C}$, the closure of a bounded Lipschitz domain in $\mathbb{K}^{n}$ and the closure of a bounded domain in $\mathbb{K}^{n}$ with the geometrical condition, explained in Definition 1.8, are shown to have the local Hölder continuity property as subsets of $\mathbb{C}^{n}$ in [Ng24, Example 4.5]. In particular, the real cube $0,1]^{n}$, the polydisk $\bar{\mathbb{B}}({0},1)^{n}$ and a convex compact subset of $\mathbb{K}^{n}$ with the nonempty interior in $\mathbb{K}^{n}$ has a uniform density in capacity in $\mathbb{C}^{n}$. In [Ng24, Example 4.5], the uniform interior sphere condition implies the geometrical condition since the union of uniform closed balls conjecture was proved by Nour, Stern and Takche [NST11]. [Ng24, Lemma 4.7-(a)] directly showed that the closure of a bounded domain in $\mathbb{C}^{n}$ with a uniform interior sphere condition implies the local Hölder continuity property with improvement in both the exponent and the order compared to the geometric condition. This direct proof is also valid since the uniform interior $r$-sphere condition implies the union of uniform closed $r/2$-balls by [NST11].
@@
$\alpha(K):=\frac{2\mu}{q+2}\quad\text{and}\quad\alpha^{\prime}(K):=\frac{4\mu^{2}}{(q+2)(2\mu+(q+2)(nq+2))}$
@@
The Hölder continuity property of a general compact subset in $\mathbb{C}^{n}$, introduced in Section 3, does not imply the local Hölder continuity property, introduced in Section 3, of the subset. However, if a compact subset in $\mathbb{C}^{n}$ is convex, then we prove that the Hölder continuity property of the subset implies the local Hölder continuity property of it with the same exponent and the order equal to the exponent. We also prove that the Hölder continuity property of a compact subset at a star center of the set implies the local Hölder continuity property of the set at the point with the same exponent and the order equal to the exponent.
@@
Let $F$ be a compact subset of $\mathbb{C}^{n}$ and let $\mu,\nu\in(0,1]$ be constants. If $F$ is convex and has $\mu$-Hölder continuity property, then $F$ has local $\mu$-Hölder continuity property of order $\mu$. If $F$ has $\nu$-Hölder continuity property at $b\in F$ and $b$ is a star center of $F$, i.e., $tF+(1-t)b\subset F$ for each $t\in[0,1]$, then $F$ has local $\nu$-Hölder continuity property of order $\nu$ at $b$.
@@
Let $\mathbb{K}$ be either $\mathbb{R}$ or $\mathbb{C}$. Then a convex compact subset of $\mathbb{K}^{n}$ with nonempty interior in $\mathbb{K}^{n}$ has local $1/2$-Hölder continuity property of order $1/2$.
@@ 2. Basic properties of extremal functions
The basic properties introduced in this section are the analogues for the extremal functions of the compact sets in $\mathbb{C}^{n}$. For the details of these basic properties, we refer to [Ah26]. Note that for subsets $E$ and $O$ of $X$ (resp. of $\mathbb{C}^{n}$), if $O$ is open in $X$ (resp. in $\mathbb{C}^{n}$), then $\sup_{O}V_{E}=\sup_{O}V_{E}^{*}$ (resp. $\sup_{O}L_{E}=\sup_{O}L_{E}^{*}$).
@@ 3.2. Local Hölder continuity property
First, assume that $F$ is convex and $F$ has $\mu$-HCP. By the HCP condition, there exists a constant $C>0$ such that
$L_{F}(z)\leq Cd_{\mathbb{C}^{n}}(z,F)^{\mu},\quad d_{\mathbb{C}^{n}}(z,F)\leq 1.$
Since $L_{F}^{*}\in\mathcal{L}$ and $L_{F}^{*}$ is locally bounded from above, for the constant
$C^{\prime}:=\max\{C,\sup_{d_{\mathbb{C}^{n}}(z,F)>1}\frac{L_{F}(z)}{d_{\mathbb{C}^{n}}(z,F)^{\mu}}\}>0,$
we have
(3.2) $L_{F}(z)\leq C^{\prime}d_{\mathbb{C}^{n}}(z,F)^{\mu},\quad z\in\mathbb{C}^{n}.$
Since $F$ is convex, for each $a\in F$ and $0<r\leq 1$, the affinely transformed set a+(Fa)/ra+(F-a)/r contains $F$. Since $F$ is compact, there exists a small constant $\epsilon\in(0,1)$ such that $\bar{\mathbb{B}}(a,1/\epsilon)$ contains $F$ for each $a\in F$. Then we have
$L_{F\cap\bar{\mathbb{B}}(a,r)}(z)$ $=L_{(a+\frac{F-a}{r\epsilon})\cap\bar{\mathbb{B}}(a,\frac{1}{\epsilon})}(a+\frac{z-a}{r\epsilon})$
$\leq L_{F}(a+\frac{z-a}{r\epsilon})$
$\leq{C^{\prime}}{d_{\mathbb{C}^{n}}(a+\frac{z-a}{r\epsilon},F)^{\mu}}$
$\leq C^{\prime}\|\frac{z-a}{r\epsilon}\|^{\mu}$
$=\frac{C^{\prime}}{\epsilon^{\mu}}\frac{\|z-a\|^{\mu}}{r^{\mu}},\quad z\in\mathbb{C}^{n},0<r\leq 1.$
Therefore, $F$ has local $\mu$-HCP of order $\mu$.
For the second part, assume that $b\in F$ is a star center of $F$ and $F$ has $\nu$-HCP at $b$. By the HCP of $F$ at $b$, there exists a constant $C_{b}>0$ such that
$L_{F}(z)\leq C_{b}\|z-b\|^{\nu},\quad\|z-b\|\leq 1.$
Since $L_{F}^{*}\in\mathcal{L}$ and $L_{F}^{*}$ is locally bounded from above, for the constant
$C_{b}^{\prime}:=\max\{C_{b},\sup_{\|z-b\|>1}\frac{L_{F}(z)}{\|z-b\|^{\nu}}\}>0,$
we have
$L_{F}(z)\leq C_{b}^{\prime}d_{\mathbb{C}^{n}}(z,F)^{\nu},\quad z\in\mathbb{C}^{n}.$
Since $b$ is a star center of $F$, for each $t\in(0,1]$, the set b+(Fb)/tb+(F-b)/t contains $F$. Since $F$ is compact, there exists $\epsilon_{b}\in(0,1)$ such that $\bar{\mathbb{B}}(b,1/\epsilon_{b})$ contains $F$. Then
$L_{F\cap\bar{\mathbb{B}}(b,r)}(z)$ $=L_{(b+\frac{F-b}{r\epsilon_{b}})\cap\bar{\mathbb{B}}(b,\frac{1}{\epsilon_{b}})}(b+\frac{z-b}{r\epsilon_{b}})$
$\leq L_{F}(b+\frac{z-b}{r\epsilon_{b}})$
$\leq\frac{C_{b}^{\prime}}{\epsilon_{b}^{\nu}}\frac{\|z-b\|^{\nu}}{r^{\nu}},\quad z\in\mathbb{C}^{n},0<r\leq 1.$
Therefore, $F$ has local $\nu$-HCP of order $\nu$ at $b$. ∎
Remark 3.3.
By the inequality  (3.2), a compact subset $F$ of $\mathbb{C}^{n}$ has HCP if and only if $L_{F}$ is Hölder continuous on $\mathbb{C}^{n}$. Suppose a compact subset $F\subset\mathbb{C}^{n}$ has $\mu$-HCP with a Hölder coefficient $C$ as
@@
$L_{F}(z)\leq 2C\max\{d_{\mathbb{C}^{n}}(z,F),d_{\mathbb{C}^{n}}(z,F)^{\mu}\},\quad z\in\mathbb{C}^{n}.$
By this inequality, we know that in the proof of Theorem 1.7, in the case of $\mu=1$, $2C$ can serve as $C^{\prime}$.
@@ 4.1. Demailly estimates
(4.3), (4.5) and [GKZ08, Remark 4.2](or [Z20, Theorem 3.2]) together give
@@ 5.3. Proof of the third theorem
[Mc34], Mcshane extension, extends a Hölder continuous function from a compact subset to $X$ with the same Hölder exponent and the same Hölder constant. Since $X$ is compact, each connected component of $X$ has a finite diameter with respect to the Riemannian distance $d_{\omega}$. Therefore, for a given bounded subset $\{\phi_{\gamma}\}_{\gamma\in\Gamma}$ in $C^{\mu^{\prime}}(K)$, each $\phi_{\gamma}$ can be extended to a function $\tilde{\phi}_{\gamma}$ in $C^{\mu^{\prime}}(X)$ such that $\{\tilde{\phi}_{\gamma}\}_{\gamma\in\Gamma}$ is bounded in $C^{\alpha}(X)$. Let $\mu^{\prime\prime}:=(\mu^{\prime})^{2}/(\mu^{\prime}+2+nq)$. The direct computation gives
$\mu^{\prime\prime}=\frac{4\mu^{2}}{(q+2)(2\mu+(q+2)(nq+2))}.$